# Quantum Rainbow Codes

### v2

@@ C.1 Maximal Subgraphs
Maximal subgraphs of type $S$ can be obtained by finding connected components of the simplex graph where vertices are considered adjacent only if the edge is of colour $c_{i}\in S$. In Algorithm 1, we build maximal subgraphs recursively from subgraphs with a smaller number of colours using a spanning tree algorithm. To calculate the spanning tree, we consider sets of subgraphs $F$ connected by a set of edges of colour $c_{0}$ see Algorithm 2.
@@
         CC,ST,Cycles = SpanningTree(G,c0,F)
         return [vertices(C) for C in CC]
@@ C.2 Rainbow Subgraphs
Rainbow subgraphs are more complex to obtain in general than maximal subgraphs. Rainbow subgraphs of type $S$ are subgraphs of a maximal subgraph of type $S$. Finding rainbow subgraphs exhaustively would require us to consider all possible combinations of vertices within each maximal subgraph of the corresponding type.
@@ C.2.1 Two-Colour Rainbow Subgraphs
In Algorithm 3, we show how to construct two-colour rainbow subgraphs of type $S=\{c_{0},c_{1}\}$. We do this by first constructing the maximal subgraphs $F$ of type $\{c_{1}\}$ using Algorithm 1. We then construct a spanning tree where subgraphs in $F$ are connected by edges of colour $c_{0}$ using Algorithm 2. This results in a set of connected components, a spanning tree and a set of unvisited edges which correspond to cycles. Unvisited edges are represented by a tuple $(f_{1},f_{2},v_{1},v_{2})$ where $(v_{1},v_{2})$ is an edge of colour $c_{0}$ joining two maximal subgraphs $f_{1},f_{2}$. Using Algorithm 4, we find the common ancestor $f$ of $f_{1}$ and $f_{2}$ and the set of vertices joining $f,f_{1}$ and $f_{2}$ which form the rainbow subgraph.
@@ C.2.2 Multi-colour Colour Rainbow Subgraphs
In this work, we construct CSS codes whose Z-checks are associated with $2$-colour subgraphs and X-checks with $D$-colour subgraphs. As a result, Z-checks can be generated efficiently, whether they are maximal (using Algorithm 1) or rainbow type (using Algorithm 3). X-checks which are of maximal type can also be generated efficiently, leaving X-checks of rainbow type $S$ where there are more than two colours in $S$.
In Algorithm 5, we show how to generate the remaining X-checks of rainbow type $S$. We first calculate the maximal subgraphs of type $S$ using Algorithm 1. Rainbow subgraphs of type $S$ are contained within a maximal subgraph of type $S$. Any X-check must also commute with the Z-checks and so be in the intersection of a maximal subgraph and the kernel of the Z-checks. The intersection of spans can be calculated using linear algebra techniques as set out in Algorithm 6.
@@
A list of colours c = [c0,..,cd] where d $>$ 1
@@ Appendix D Coloured Logical Paulis
In this Appendix, we show how to generate the coloured Paulis as defined in Section IV.B of [10]. We set out a method to calculate coloured logical Z operators. Swapping Z-checks and logicals with X-checks and logicals results in a method for logical X operators. One of the inputs to the algorithm is a generating set of logical Z operators which are not necessarily coloured logicals. These can be generated, for instance, by using the method in Section 10.5.7 of [39].

### v3

@@ C.1 Maximal Subgraphs
Maximal subgraphs of type $S$ can be obtained by finding connected components of the simplex graph where vertices are considered adjacent only if the edge is of colour $c_{i}\in S$. In Algorithm˜1, we build maximal subgraphs recursively from subgraphs with a smaller number of colours using a spanning tree algorithm. To calculate the spanning tree, we consider sets of subgraphs $F$ connected by a set of edges of colour $c_{0}$ - see Algorithm˜2. The algorithm for computing maximal subgraphs has complexity linear in the number of edges and vertices of the simplex graph.
@@
   $C$,ST,Cycles = SpanningTree(G,c0,F)
   return [vertices(C) for C in $C$]
@@ C.2 Rainbow Subgraphs
Rainbow subgraphs are more complex to obtain in general than maximal subgraphs. Rainbow subgraphs of type $S$ are subgraphs of a maximal subgraph of type $S$. Finding rainbow subgraphs exhaustively would require us to consider all possible combinations of vertices within each maximal subgraph of the corresponding type, and is worst-case exponential complexity in the number of vertices in the maximal subgraphs. We demonstrate below efficient algorithms for finding rainbow subgraphs in the special case for rainbow codes where the Z-type checks are two-colour rainbow subgraphs.
@@ C.2.1 Two-Colour Rainbow Subgraphs
In Algorithm˜3, we show how to construct two-colour rainbow subgraphs of type $S=\{c_{0},c_{1}\}$. We do this by first constructing the maximal subgraphs $F$ of type $\{c_{1}\}$ using Algorithm˜1. We then construct a spanning tree where subgraphs in $F$ are connected by edges of colour $c_{0}$ using Algorithm˜2. This results in a set of connected components, a spanning tree and a set of unvisited edges which correspond to cycles. Unvisited edges are represented by a tuple $(f_{1},f_{2},v_{1},v_{2})$ where $(v_{1},v_{2})$ is an edge of colour $c_{0}$ joining two maximal subgraphs $f_{1},f_{2}$. Using Algorithm˜4, we find the common ancestor $f$ of $f_{1}$ and $f_{2}$ and the set of vertices joining $f,f_{1}$ and $f_{2}$ which form the rainbow subgraph. The complexity of this algorithm is linear in the number of vertices and edges of the simplex graph, but also depends on the number of vertices in the rainbow subgraphs.
@@ C.2.2 Multi-colour Colour Rainbow Subgraphs
In this work, we construct CSS codes whose Z-checks are associated with $2$-colour subgraphs and X-checks with $D$-colour subgraphs. As a result, Z-checks can be generated efficiently, whether they are maximal (using Algorithm˜1) or rainbow type (using Algorithm˜3). X-checks which are of maximal type can also be generated efficiently, leaving X-checks of rainbow type $S$ where there are more than two colours in $S$.
In Algorithm˜5, we show how to generate the remaining X-checks of rainbow type $S$. We first calculate the maximal subgraphs of type $S$ using Algorithm˜1. Rainbow subgraphs of type $S$ are contained within a maximal subgraph of type $S$. Any X-check must also commute with the Z-checks and so be in the intersection of a maximal subgraph and the kernel of the Z-checks. The intersection of spans can be calculated using linear algebra techniques as set out in Algorithm˜6. The kernel calculation dominates the complexity of this algorithm and accordingly the overall complexity is polynomial in the number of vertices in the simplex graph.
@@
A list of colours c = [c0,..,cd] where d >> 1
@@ Appendix D Coloured Logical Paulis
In this Appendix, we show how to generate the coloured Paulis as defined in Section IV.B of [vuillot2022quantum]. We set out a method to calculate coloured logical Z operators. Swapping Z-checks and logicals with X-checks and logicals results in a method for logical X operators. One of the inputs to the algorithm is a generating set of logical Z operators which are not necessarily coloured logicals. These can be generated, for instance, by using the method in Section 10.5.7 of [nielsen_quantum_2010].